# The anonymization problem in social networks

### v1

@@ Appendix A Recompute gap
This appendix accompanies Section 5.2 of the full paper. Here, we compare the performance and runtimes achieved by the anonymization algorithms when using different recompute gaps. Figure 5 shows the results for different algorithms when using different recompute gaps. The value of 1 implies that values are recomputed after each edge deletion, $|E|/100$ equals the default setting. The figure overall shows that using a smaller recompute gap similar results are obtained. Turning to the runtimes in Figure 6 we see that using a smaller recompute gap does result in larger runtimes.
@@ Appendix B Anonymization process
This appendix contains supplementary figures accompanying Section 6.3 up to 6.5 of the full paper. Results shows the full anonymization process, i.e., up to deleting all edges. Figure 7 shows results for all anonymity measures for all networks not shown in Figure 1 of the full paper. Figure 8 shows this process for the count measure.
@@
Figure 8: Anonymization process for all networks in Table 2 of the paper and measures (color and linestyle) for count.
@@ Appendix C Properties and anonymization
This appendix accompanies Section 6.3 and 6.4 of the full paper and contains results on correlations between network properties and the effectiveness of the anonymization algorithms on the three variants of the anonymization problem. Table 3 denotes the Pearson correlation and p-value. For combinations deemed significant, Figure 9 shows the values for all networks.

### v2

@@ Appendix A Recompute gap
This appendix accompanies Section 5.2 of the full paper. Here, we compare the performance and runtimes achieved by the anonymization algorithms when using different recompute gaps. Figure 7 shows the results for different algorithms on a subset of the network datasets when using different recompute gaps. The recompute gap value of 1 implies that values are recomputed after each edge deletion, $|E|/100$ equals the default setting. The figure overall shows that using a smaller recompute gap similar results are obtained. Turning to the runtimes in Figure 8 we see that using a smaller recompute gap does result in larger runtimes.
@@ Appendix B Anonymization process
This appendix contains supplementary figures accompanying Section 6.3. Results shows the full anonymization process, i.e., up to deleting all edges. Figure 9 shows results for all anonymity measures for all networks not shown in Figure 2 of the full paper.
@@ Appendix C Properties and anonymization
This appendix accompanies Section 6.4 of the full paper and contains results on correlations between network properties and the effectiveness of the anonymization algorithms on the three variants of the anonymization problem. Table 3 denotes the Pearson correlation and p-value. For combinations deemed significant, Figure 10 shows the values for all networks.
@@
Appendix D Deleting edges
This appendix accompanies Sections 6.6 of the full paper. Each figure shows how a network property or performance on a downstream task changes when deleting more edges, using the five anonymization algorithms.
Figure 11: Change in network properties when deleting edges. For each network (rows) and property (columns) the results show how data utility changes when deleting more edges (horizontal axis) using each of the five anonymization algorithms (color).
Appendix E Pareto fronts
This appendix accompanies Section 6.7 of the full paper. Figure 12 shows for each network the Pareto optimal solutions found by the algorithms, for each of the utility metrics.
Figure 12: Pareto fronts in terms of uniqueness (horizontal axis) and utility (vertical axis). For each network (rows), the results show for each property (columns) the Pareto optimal points found by each algorithm (color). Lines show one standard deviation.
Appendix F Runtimes
This appendix accompanies Section 6.8 of the full paper. Each figure shows both the full runtime, and time spent on the anonymization algorithm for a given network and anonymization algorithm.
Figure 13: Runtime per iteration. Each figure shows for each combination of network (rows) and algorithm (columns and color) the runtime per iteration. The color corresponding to the algorithm indicates the full runtime, the black dots the time spent on the anonymization algorithm only.